# Safe and Responsible Large Language Model Development

### v2

@@ 6 Conclusion
Appendix A Annotation Guidelines for CMD
@@
Miscellaneous: Ageism, Disabled, Diseased, Feminist, Physical Appearance, Socio-economic Status.
Following are the examples for annotation guidelines.
@@
| Category | Fleiss Kappa Score | Interpretation (Agreement) |
| Toxicity | 0.75 | Substantial |
| Harm | 0.72 | Substantial |
| Bias | 0.78 | Substantial |
| (-ve) Sentiment | 0.70 | Substantial |
| Overall | 0.69 | Substantial |
In this annotation process, we evaluated texts for unsafe content, focusing on bias, toxicity, negative sentiment, and harm. The Fleiss’ Kappa scores, a measure of inter-annotator agreement, were thoroughly calculated for each category. For toxicity, we observed a Fleiss’ Kappa score of 0.75, indicating strong agreement. The harm category showed a substantial consensus with a score of 0.72. Bias identification and removal also achieved a high degree of agreement, with a score of 0.78. Negative sentiment analysis resulted in a good agreement score of 0.70. These scores represent the annotators’ uniform understanding and application of the comprehensive guidelines established.
Overall, the aggregate Fleiss’ Kappa score across all categories was approximately 0.69, as detailed in 10. This score signifies a strong consensus among our diverse team of 15 annotators, comprising 5 experts and their paired students, on the reliability and consistency of the annotations. The process involved evaluating a balanced set of 20,000 records, aiming for diversity and computational efficiency in LLM fine-tuning. Our approach, validated by the consistent Fleiss’ Kappa scores, ensures the methodological soundness of our dataset, underlining the effectiveness of our annotation guidelines and the annotators’ adherence to these principles.
@@
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - Medium, Target - Millennials, Annotation - Unsafe
@@
Details: Bias - Yes, Toxicity - No, Sentiment - Negative, Harm - Medium, Target - Women, Annotation - Unsafe
@@
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - High, Target - Non-English Speakers, Annotation - Unsafe
@@ B.1 Data Schema
B.2 Data Analysis
| Statistic | char_length | word_length |
| Count | 16614.000 | 16614.000 |
| Mean | 370.336 | 68.277 |
| Std | 451.732 | 80.136 |
| Min | 7.000 | 1.000 |
| 25% | 110.000 | 22.000 |
| 50% | 238.000 | 44.000 |
| 75% | 491.000 | 90.000 |
| Max | 5000.000 | 1108.000 |
Table 11 shows the text length statistics indicate moderate average lengths but with considerable variability, suggesting the dataset contains a diverse array of text sizes, which is advantageous for creating versatile NLP models. The labels distribution is in Appendix12.
| Label | Category (Count) |
| BIAS | No (14227), Yes (5772) |
| TOXICITY | No (12040), Mild (5293), High (2666) |
| SENTIMENT | Negative (9028), Neutral (8370), Positive (2601) |
| HARM | Low (14151), Medium (3932), High (1915) |
| ANNOT | Unsafe (10,359) , Benign (9,640) |
Figure 6 Flesch Reading Ease scores imply that positive sentiments are found in more straightforward texts, while negative and neutral sentiments appear in more complex ones, potentially affecting the ease of automated readability assessment.
| Statistic | Value |
| Count | 8307.000000 |
| Mean | 0.087286 |
| Std | 0.148290 |
| Min | 0.000000 |
| 25% | 0.028708 |
| 50% | 0.052493 |
| 75% | 0.084249 |
| Max | 1.000000 |
Table 13 demonstrates the benign analysis of text alterations post-de-biasing, revealing a notable variation in the extent of change. This highlights the complexity of the debiasing process and the need for careful consideration in maintaining the original text content.
@@ Appendix C QLoRA Details
QLoRA (Quantization with Low-Rank Adapters) Dettmers et al. [2023] is an efficient fine-tuning technique used in the context of LLaMA-2. LLaMA-2 is a powerful pre-trained model that can be fine-tuned for various natural language understanding tasks. Fine-tuning involves adapting a pre-trained model to a specific downstream task by training it on task-specific data. It helps the model specialize and perform well on specific tasks. QLoRA is the key technique used during fine-tuning. QLoRA quantizes the pre-trained LLaMA-2 model to use fewer bits for its weights, specifically reducing the precision of the model’s weights to 4 bits. After quantization, QLoRA attaches small “Low-Rank Adapters” to the model, which are fine-tuned on the task-specific data.
@@
Packages We utilize models through HuggingFace Transformers Wolf et al. [2020], and utilize the Trainer from this package to train our models. In addition, we use integrated bitsandbytes for Parameter Efficient Finetuning.
@@ Appendix D Carbon Footprint Estimation
Quantifying the carbon footprint Dodge et al. [2022] of machine learning (ML) model training involves calculating the energy consumption during the training process and converting this figure into carbon dioxide equivalent (CO2e) emissions. This calculation is informed by the power usage of the computing hardware (GPUs and CPUs) and the carbon intensity of the electrical supply, which can vary significantly by geographic location and energy source. For our estimations, we adopt a global average carbon intensity of about 0.4 kgCO2e per kilowatt-hour (kWh) as a standard metric.
@@
The process for calculating energy consumption in kWh encompasses the following steps:
- Total Power Consumption: Summation of the GPUs’ and CPUs’ power usage in Watts.
@@
- Calculated Energy Use (kWh): Multiplication of the total power usage (in kW) by the session duration in hours, yielding the total energy consumption.
Carbon Emission Estimation:
@@ E.1 Classifiers
OpenAI Moderation API OpenAI [2024]employs a GPT-based, multi-label classifier specifically fine-tuned to evaluate whether text breaches one of eleven content safety categories. These categories encompass hate, harassment, self-harm, sexual content involving minors, violence, and more. For each category, the endpoint delivers a probability score and a binary label, alongside an overall binary label indicating the content’s safety status.
@@
The Perspective API API [2024]serves as a tool for online platforms and publishers to identify and remove harmful or offensive content, particularly within comments and discussions. Leveraging machine learning models, it analyzes text to provide probability scores indicating the likelihood of it being perceived as harmful. The risk categories assessed by Perspective API include toxicity, severe toxicity, identity attack, insult, profanity, and threat.
@@
Style Classifier Smith et al. [2020]: The style classifier rom ParlAI identifies specific attributes or patterns in text, such as sentiment, formality, or writing style, it is a pre-trained model. In the context of CLEN (Content-Length Entropy Normalization), is a metric that measures the diversity of content length in generated text. Higher CLEN values indicate more varied sentence lengths, which can impact readability and style coherence.
Our fine-tuned classifiers
Bias Classifier: This classifier has been fine-tuned on a dataset labeled for bias (biased, non-biased), using the advanced capabilities of BERT-Large for thorough detection and classification of bias. The model underwent fine-tuning on an extensive dataset comprising 3.7 million records. This classifier is publicly available for use.
Toxicity Classifier We use the default ToxiGen classifier tuned on RoBERTa to measure the toxicity of generations of each of the LLMs.
Sentiment Classifier The Sentiment Classifier was extensively fine-tuned on our dataset with sentiment data labels (positive, neutral, negative) using BERT-Large. This fine-tuning enhances its ability to perform nuanced sentiment analysis across various text inputs. This classifier is publicly available for use.
Harm Classifier: For the Harm Classifier, we fine-tuned it using our training dataset, which includes specific labels for harm (low, medium, high) with DeBERTa-Large. This process ensures the accurate identification of harmful content and behaviors.This classifier is publicly available for use.
Unsafe Classifier: We meticulously fine-tuned the RoBERTa-Large model using our training dataset to build the Unsafe Classifier. This dataset’s class labels are human annotations (ANNOT.) that label content as either unsafe or benign.
E.2 Metrics
Fairness Metrics
We adopt the fairness metrics from Steroeset Nadeem et al. [2021], which are:
Language Modeling Score (LMS): The LMS assesses the model’s baseline language modeling ability. A perfect score of 100 indicates that the model correctly associates and understands every word or phrase in the given context, providing meaningful responses.
Stereotype Score (SS): The SS measures the model’s tendency towards stereotype or anti-stereotype terms. A score of 50 signifies a neutral stance, indicating no inherent bias towards stereotypical terms. Scores deviating from 50 indicate the model’s inclination towards either stereotype or anti-stereotype terms.
Idealized Context Association Test (ICAT): The ICAT evaluates a model’s performance by combining LMS and SS. These metrics collectively provide insights into the model’s language understanding abilities and its inclination towards stereotypes or anti-stereotypes. The ICAT score is calculated by combining the LMS and SS. An ideal model achieving an LMS of 100 and an SS of 50 would attain an ICAT score of 100, reflecting impeccable language modeling abilities and a neutral stance towards stereotypes. Conversely, a model demonstrating complete stereotypical bias, with an LMS of 0 and an SS of either 0 or 100, would receive an ICAT score of 0, indicating poor language modeling abilities and a strong bias towards stereotypes.
The Partial Gen Bias metric Smith et al. [2022] in the HolisticBias context refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model’s ability to produce content that exhibits bias without being overtly toxic or offensive. The goal is to capture nuanced biases present in language generation, which may not be immediately apparent but still contribute to societal prejudices.
CLEN (Content-Length Entropy Normalization) Miller et al. [2017] is a metric employed to gauge the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contibutes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
T-Tests
A t-test is a statistical method used to determine if there is a significant difference between the means of two groups or to compare the mean of a single group against a known standard. It helps in understanding if any observed differences are due to chance or if they reflect actual differences in the populations Kim [2015].
One-Sample T-Test Ross and Willson [2017]: To measure the effect of safety interventions on the texts before and after the implementation of these measures, we employ the one-sample t-test. By comparing the mean value of a specific characteristic of the texts (e.g., the proportion of texts classified as safe) before and after the intervention against a specific benchmark or expected value, we statistically evaluate the effectiveness of our safety intervention in this work. The formula for the one-sample t-test is:
where:
is the sample mean,
is the population mean,
is the sample standard deviation,
is the sample size.
E.3 Safety Prompts

### v3

@@ Acknowledgements
Appendix A Annotation Guidelines
@@
Miscellaneous: Ageism, Disabled, Diseased, Feminist, Physical Appearance, Socio-economic Status.Following are the examples for annotation guidelines.
@@
Table 9: Summary of Fleiss’ Kappa Scores and Interpretations Category Fleiss Kappa Score Interpretation (Agreement) Toxicity 0.75 Substantial Harm 0.72 Substantial Bias 0.78 Substantial (-ve) Sentiment 0.70 Substantial Benign variation 0.78 Substantial The Fleiss’ Kappa scores, a measure of inter-annotator agreement, were thoroughly calculated for each category. For toxicity, we observed a Fleiss’ Kappa score of 0.75, indicating strong agreement. The harm category showed a substantial consensus with a score of 0.72. Bias identification and removal also achieved a high degree of agreement, with a score of 0.78. Negative sentiment analysis resulted in a good agreement score of 0.70. The benign generation score is calculated to be 0.78.
@@
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - Medium, Target - Millennials
@@
Details: Bias - Yes, Toxicity - No, Sentiment - Negative, Harm - Medium, Target - Women
@@
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - High, Target - Non-English Speakers,
@@ B.1 Data Schema
Listing 1: Dataset Schema {"$schema": "http://json-schema.org/draft-07/schema#","title": "Dataset","description": "A␣dataset␣schema␣for␣analyzing␣sentences␣for␣bias,␣toxicity,␣sentiment,␣harm,␣and␣a␣benign␣variation␣for␣a␣unsafe␣text.","type": "object","properties": {"Original␣Sentence": {"type": "string","description": "The␣text␣of␣the␣original␣sentence␣that␣may␣be␣unsafe"},"BIAS": {"type": "string","description": "Indicates␣if␣the␣sentence␣is␣biased.","enum": ["Yes", "No"]},"TOXICITY": {"type": "string","description": "Level␣of␣toxicity␣of␣the␣sentence.","enum": ["No", "Mild", "High"]},"SENTIMENT": {"type": "string","description": "Sentiment␣of␣the␣sentence.","enum": ["Negative", "Neutral", "Positive"]},"HARM": {"type": "string","description": "Level␣of␣harm␣of␣the␣sentence.","enum": ["Low", "Medium", "High"]},"DEMOGRAPHIC␣TARGETING": {"type": "string","description": "Indicates␣if␣specific␣demographics␣are␣targeted.","enum": ["None", "Specific␣Demographics"]},"WORDS␣OR␣PHRASES": {"type": "string","description": "List␣of␣biased␣words␣or␣phrases␣identified␣in␣the␣sentence."},"Benign": {"type": "string","description": "The␣benign␣version␣of␣the␣text."},"required": ["Original␣Sentence", "BIAS", "TOXICITY", "SENTIMENT", "HARM", "Benign"]}B.2 Data Analysis
Table 10: Descriptive Statistics for Text Length Statistic char_length word_length Count 16614.000 16614.000 Mean 370.336 68.277 Std 451.732 80.136 Min 7.000 1.000 25% 110.000 22.000 50% 238.000 44.000 75% 491.000 90.000 Max 5000.000 1108.000 Table 10 shows the text length statistics indicate moderate average lengths but with considerable variability, suggesting the dataset contains a diverse array of text sizes, which is advantageous for creating versatile NLP models. The labels distribution is in Appendix11.
Label Category (Count) BIAS No (14227), Yes (5772) TOXICITY No (12040), Mild (5293), High (2666) SENTIMENT Negative (9028), Neutral (8370), Positive (2601) HARM Low (14151), Medium (3932), High (1915) Table 11: Summary of Label Distributions
Figure 6: Flesch Reading Ease Scores and its Relationship with Sentiment Label Figure 6 Flesch Reading Ease scores imply that positive sentiments are found in more straightforward texts, while negative and neutral sentiments appear in more complex ones, potentially affecting the ease of automated readability assessment.
Statistic Value Count 8307.000000 Mean 0.087286 Std 0.148290 Min 0.000000 25% 0.028708 50% 0.052493 75% 0.084249 Max 1.000000 Table 12: Analyzing the Impact of Benign Table 12 demonstrates the benign analysis of text alterations post-de-biasing, revealing a notable variation in the extent of change. This highlights the complexity of the debiasing process and the need for careful consideration in maintaining the original text content.
@@ Appendix C QLoRA Details
QLoRA (Quantization with Low-Rank Adapters) [30] is an efficient fine-tuning technique used in the context of LLaMA-2. LLaMA-2 is a powerful pre-trained model that can be fine-tuned for various natural language understanding tasks. Fine-tuning involves adapting a pre-trained model to a specific downstream task by training it on task-specific data. It helps the model specialize and perform well on specific tasks. QLoRA is the key technique used during fine-tuning. QLoRA quantizes the pre-trained LLaMA-2 model to use fewer bits for its weights, specifically reducing the precision of the model’s weights to 4 bits. After quantization, QLoRA attaches small “Low-Rank Adapters” to the model, which are fine-tuned on the task-specific data.
@@
Packages We utilize models through HuggingFace Transformers [46], and utilize the Trainer from this package to train our models. In addition, we use integrated bitsandbytes for Parameter Efficient Finetuning.
@@ Appendix D Carbon Footprint Estimation
Quantifying the carbon footprint [31] of machine learning (ML) model training involves calculating the energy consumption during the training process and converting this figure into carbon dioxide equivalent (CO2e) emissions. This calculation is informed by the power usage of the computing hardware (GPUs and CPUs) and the carbon intensity of the electrical supply, which can vary significantly by geographic location and energy source. For our estimations, we adopt a global average carbon intensity of about 0.4 kgCO2e per kilowatt-hour (kWh) as a standard metric.
@@
The process for calculating energy consumption in kWh encompasses the following steps: - Total Power Consumption: Summation of the GPUs’ and CPUs’ power usage in Watts.
@@
- Calculated Energy Use (kWh): Multiplication of the total power usage (in kW) by the session duration in hours, yielding the total energy consumption.Carbon Emission Estimation:
@@ E.1 Classifiers
OpenAI Moderation API [40]employs a GPT-based, multi-label classifier specifically fine-tuned to evaluate whether text breaches one of eleven content safety categories. These categories encompass hate, harassment, self-harm, sexual content involving minors, violence, and more. For each category, the endpoint delivers a probability score and a binary label, alongside an overall binary label indicating the content’s safety status.
@@
The Perspective API [39]serves as a tool for online platforms and publishers to identify and remove harmful or offensive content, particularly within comments and discussions. Leveraging machine learning models, it analyzes text to provide probability scores indicating the likelihood of it being perceived as harmful. The risk categories assessed by Perspective API include toxicity, severe toxicity, identity attack, insult, profanity, and threat.
@@
Style Classifier [45]: The style classifier rom ParlAI identifies specific attributes or patterns in text, such as sentiment, formality, or writing style, it is a pre-trained model. In the context of CLEN (Content-Length Entropy Normalization), is a metric that measures the diversity of content length in generated text. Higher CLEN values indicate more varied sentence lengths, which can impact readability and style coherence.
E.2 Safety Prompts